i don't understand how to add 5 3/5 + 2 2/3
Answer by Alan3354(69443)

i don't understand how to add 5 3/5 + 2 2/3
-----------------
5 + 2 = 7
3/5 + 2/3 = 9/15 + 10/15 = 19/15 = 1 4/15
7 + 1 4/15 = 8 4/15
------
Or,
5 3/5 + 2 2/3 = 28/5 + 8/3
= 84/15 + 40/15 = 124/15
----
Mixed numbers are just a nuisance anyway.